Lucky Number: Question 4. (15 points) An individual who has automobile insurance from a certain company is randomly selected. Let Y be the number of moving violations for which the individual was cited during the last 3 years. The pmf ofY is Page 4 of 8 p(y) 0.25 0.10 0.05 (1). Find a, such that p(y) is a p.mf. (2).Write out the c.d.f of Y completely. (3).Suppose an individual with Y violations incurs a surcharge of S(+100y2). Calculate the expected amount of the surcharge.
Answer #1

Ans:

1)

P(y=0)=1-0.25-0.1-0.05

=1-0.4=0.6

2)

y p(y) cdf=P(Y<=y)
0 0.6 0.6
1 0.25 0.85
2 0.1 0.95
3 0.05 1
Total 1

3)

E(y)=0*0.6+1*0.25+2*0.1+3*0.05

=0+0.25+0.2+0.15=0.6

E(y^2)=0^2*0.6+1^2*0.25+2^2*0.1+3^2*0.05=1.1

Expected value of surcharge=E(Y+100Y^2)

   =E(Y)+100*E(Y^2)

=0.6+100*1.1

=0.6+110=110.6